integral of e^(-x^2)
1
Expert's answer
2020-10-21T15:18:04-0400

"Integral \ of \ {e^{-x^2}}"


Firstly, set the integral to be equal to I

"I = \int{e^{-x^2}dx}"


Squaring both sides,

"I^2 = {(\int{e^{-x^2}dx})}^2 =(\int{e^{-x^2}dx})(\int{e^{-x^2}dx})"


Since x is just a dummy variable, we can write,

"I^2=(\int{e^{-x^2}dx})(\int{e^{-y^2}dy})"


Combining, we have,

"I^2=\int^{\infin}_{-\infin}(\int^{\infin}_{-\infin}{e^{-(x^2+y^2)} \ dx)dy}"


"I^2=\int^{\infin}_{-\infin}\int^{\infin}_{-\infin}{e^{-(x^2+y^2)} \ dxdy}"


Converting from Cartesian Coordinates (x,y) to Polar Coordinates (r,θ)

"(r^2 =x^2 + y^2)"

"(""dxdy =dA = rdrd\,\theta)"


"\therefore I^2 =\iint{e^{-r^2} rdrd\,\theta}"


"I^2 =\int^{2\pi}_{0}(\int^{\infin}_{0}{e^{-r^2} r\,dr)d\,\theta}"


"I^2 =\int^{2\pi}_{0}(\int^{\infin}_{0}{re^{-r^2} dr)d\,\theta}"


"I^2 =2\pi(\int^{\infin}_{0}{re^{-r^2}dr})"


"I^2 =2\pi\int^{\infin}_{0}{re^{-r^2} dr}"


Using u substitution,

"u = r^2"


"\dfrac{du}{dr} = 2r"


"du = 2r\,dr"



"\therefore I^2 =2\pi\int^{\infin}_{u =0}{re^{-u}dr}"


Multiplying both sides of the equation by 2,

"2I^2 =2\pi\int^{\infin}_{u =0}{2re^{-u}dr}"


But du = 2rdr

"\therefore 2I^2 =2\pi\int^{\infin}_{u =0}{e^{-u}du}"


"2I^2 =2\pi[e^{0}-e^{-\infin}]"


"2I^2 =2\pi(1-0)"


"2I^2 =2\pi"


"I^2 =\pi"


"I =\sqrt\pi"


"\therefore \ The \ integral \ of \ e^{−x^2} \ is \ \sqrt{\pi}"